# The disappearance of the blue and luminous progenitor of the Type IIn SN 2010jl

### v1

@@ 1 Introduction
Type IIn supernovae (hereafter SNe IIn) is an interesting and important subclass of core-collapse SNe. They are characterized by strong and narrow (a few hundreds to about 2000 km s) emission lines, sometimes with broad bases, in the spectra (Schlegel, 1990). These features indicate dense circumstellar material (CSM) surrounding the SN, which requires enhanced mass loss (10–10 /years ) of the progenitor stars within a short timescale prior to the deaths. Therefore, SNe IIn are an effective tool to investigate the mass loss of massive stars at their final evolutionary stage.
@@ 3 SN progenitor
Near the SN position (Fig. 1), we detect a single point source on the lower-resolution pre-explosion WFPC2 images (hereafter s0) and two spatially resolved sources on the higher-resolution late-time WFC3 images (hereafter s1 and s2). The measured magnitudes of s0, s1 and s2 are listed in Table 1 and shown in Fig. 2. While s1 continues to fade on the late-time images, s2 has almost constant magnitudes in both bands. Therefore, we identify s1 as SN 2010jl and s2 as a nearby field object. The source identification is consistent with Fox et al. (2017).
@@
Next we try to decompose the spectral energy distribution (SED) of s0 to isolate the SN progenitor and derive its stellar parameters. We use a blackbody to simulate the SED of the SN progenitor; as we shall show in the next section, s2 is likely a star cluster so we adopt the BPASS v2.2 population models including interacting binaries (Eldridge et al., 2017) to model its SED. We assume that the SED of s0 is the sum of both components; we argue that this is a reasonable assumption since s2 and the SN progenitor are so close to each other while the other sources in this field are too far away to contaminate s0 significantly. We only consider Galactic extinction since the host-galaxy extinction is negligible (Smith et al., 2011). We use the pysynphot package (STScI Development Team, 2013) to synthesize model magnitudes and then compare them with the observed ones of s0 and s2. The Markov Chain Monte Carlo (MCMC) technique is used (Foreman-Mackey et al., 2013) to search for the best-fitting parameters. We find that the progenitor has an effective temperature of log()=4.26 and a bolometric luminosity log()=6.52; s2 has an initial mass of 10 and a very young age of only 5 Myr. The best-fitting SEDs are displayed in Fig 3. Figure 3 also shows the position of SN2010jl’s progenitor on the Hertzsprung-Russell (HR) diagram. This analysis shows that the progenitor of SN 2010jl is a very blue and extremely luminous star that is even above the Humphreys-Davidson (HD) limit (Humphreys & Davidson, 1994).
@@ 4 SN environment
In this section we investigate the surrounding stellar populations in the environment of SN 2010jl. An environmental analysis can provide useful insights into the SN progenitor, since most massive stars form in groups and stars within each group have very similar ages and chemical abundances. We select point sources detected on the WFC3 images with the following criteria:
@@
Additionally, a few field objects, including s2, are also much brighter than the fitted stellar isochrones; they might be star clusters, binary merger products, and/or multiple stellar systems. Assuming s2 to be a star cluster, the derived age in Section 3 is consistent with that for the younger resolved stellar population.
@@ 5 Discussion
In 2018 and 2020, SN 2010jl has faded sufficiently and the composite brightness of s1 and s2 is below the pre-explosion brightness of s0 at the near-ultraviolet wavelengths (Figs. 2 and 3). This indicates that pre-explosion source at the SN position is not a stellar cluster or a field star but rather the SN progenitor. It is likely that the SN progenitor has now disappeared. An alternative scenario is that the progenitor was in outburst before explosion and returned to a quiescent phase at late times. It is also possible that the survived progenitor is obscured by dust which is newly formed in the SN explosion (Smith et al., 2012; Maeda et al., 2013; Gall et al., 2014).
@@ acknowledgments
This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences, Grant No. XDB0550300. ZXN acknowledges support from the NSFC through grant No. 12303039, and NCS’s research is funded by the NSFC grants No. 12303051 and No. 12261141690. JFL acknowledges support from the NSFC through grants No. 11988101 and No. 11933004 and from the New Cornerstone Science Foundation through the New Cornerstone Investigator Program and the XPLORER PRIZE. This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. These observations are associated with programs 8645, 14149, 14668, 15166, and 16179. This work made use of v2.2.1 of the Binary Population and Spectral Synthesis (BPASS) models as described in Eldridge et al. (2017) and Stanway & Eldridge (2018).

### v2

@@ 1 Introduction
Type IIn supernovae (hereafter SNe IIn) is an interesting and important subclass of core-collapse SNe. They are characterized by strong and narrow (a few hundreds to about 2000 km s-1) emission lines, sometimes with broad bases, in the spectra (Schlegel, 1990). These features indicate dense circumstellar material (CSM) surrounding the SN, which requires enhanced mass loss (10-3–100 /years ) of the progenitor stars within a short timescale prior to the deaths. Therefore, SNe IIn are an effective tool to investigate the mass loss of massive stars at their final evolutionary stage.
@@ 3 SN progenitor
Near the SN position (Fig. 1), we detect a single point source on the lower-resolution pre-explosion WFPC2 images (hereafter s0) and two spatially resolved sources on the higher-resolution late-time WFC3 images (hereafter s1 and s2). The measured magnitudes of s0, s1 and s2 are listed in Table 1 and shown in Fig. 2. While s1 continues to fade on the late-time images, s2 has almost constant magnitudes in both bands. Therefore, we identify s1 as SN 2010jl and s2 as a nearby field object. We note that we didn’t run dolphot with the mode of warmstart, namely, all reported sources were detected and selected on its own. The source identification is consistent with Fox et al. (2017).
@@
We also carefully checked quality parameters calculated by dolphot. We find s0 has sharpness of 0.257 and roundness of 0.154, for comparison, s2 has sharpness of 0.747, 0.344, 0.345, 0.293 from 2015 to 2020, and roundness of 1.170, 0.561, 0.496 , 0.377; s1 has sharpness of 0.262, 0.086, 0.192, 0.302, and roundness of 0.185, 0.198 , 0.419 , 0.491. This shows s0 can’t be recognized as a blend object simply based on results of WFPC2 photometry, since the offset between its two components is less than one WFPC2 pixel.
@@
Next we try to decompose the spectral energy distribution (SED) of s0 to isolate the SN progenitor and derive its stellar parameters. We use a blackbody to simulate the SED of the SN progenitor; as we shall show in the next section, s2 is likely a star cluster so we adopt the BPASS v2.2 population models including interacting binaries (Eldridge et al., 2017) to model its SED. We assume that the SED of s0 is the sum of both components; we argue that this is a reasonable assumption since s2 and the SN progenitor are so close to each other while the other sources in this field are too far away to contaminate s0 significantly. We only consider Galactic extinction since the host-galaxy extinction is negligible (Smith et al., 2011). We use the pysynphot package (STScI Development Team, 2013) to synthesize model magnitudes and then compare them with the observed ones of s0 and s2. The Markov Chain Monte Carlo (MCMC) technique is used (Foreman-Mackey et al., 2013) to search for the best-fitting parameters. We find that the progenitor has an effective temperature of log()=4.26 and a bolometric luminosity log()=6.52; s2 has an initial mass of 104 and a very young age of only 5 Myr. The best-fitting SEDs are displayed in Fig 3. Figure 3 also shows the position of SN2010jl’s progenitor on the Hertzsprung-Russell (HR) diagram. This analysis shows that the progenitor of SN 2010jl is a very blue and extremely luminous star that is even above the Humphreys-Davidson (HD) limit (Humphreys & Davidson, 1994).
@@ 4 SN environment
In this section we investigate the surrounding stellar populations in the environment of SN 2010jl. An environmental analysis can provide useful insights into the SN progenitor, since most massive stars form in groups and stars within each group have very similar ages and chemical abundances. We performed photometry with dolphot on all WFC3 images together, so that the derived photometry in each band is the inverse variance-weighted average of those on the individual frames, and then we selected point sources from the combined photometry catalog with the following criteria:
@@
Additionally, a few field objects, including s2, are also much brighter than the fitted stellar isochrones; they might be star clusters, binary merger products, multiple stellar systems, and/or LBV-like stars in outburst. Assuming s2 to be a star cluster, the derived age in Section 3 is consistent with that for the younger resolved stellar population.
@@ 5 Discussion
In 2018 and 2020, SN 2010jl has faded sufficiently and the composite brightness of s1 and s2 is below the pre-explosion brightness of s0 at the near-ultraviolet wavelengths (Figs. 2 and 3). This indicates that pre-explosion source at the SN position is not a stellar cluster or a field star but rather the SN progenitor. It is likely that the SN progenitor has now disappeared. We can’t rule out the possibility that SN 2010ji is not a terminal explosion and the pre-explosion counterpart is still alive; in this case, the star was in outburst with higher luminosity when observed by HST in 2001 and has now returned to a fainter quiescent state at late times. It is also possible the brightness decline is (partly) due to the newly formed dust (Smith et al., 2012; Maeda et al., 2013; Gall et al., 2014).
@@ acknowledgments
This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences, Grant No. XDB0550300. ZXN acknowledges support from the NSFC through grant No. 12303039, and NCS’s research is funded by the NSFC grants No. 12303051 and No. 12261141690. JFL acknowledges support from the NSFC through grants No. 11988101 and No. 11933004 and from the New Cornerstone Science Foundation through the New Cornerstone Investigator Program and the XPLORER PRIZE. This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. These observations are associated with programs 8645, 14149, 14668, 15166, and 16179, they were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/nrbd-qp80 (catalog DOI: V). This work made use of v2.2.1 of the Binary Population and Spectral Synthesis (BPASS) models as described in Eldridge et al. (2017) and Stanway & Eldridge (2018).